# Shree Pannalal Patel Marg, Ahmedabad - 380009, Gujarat, India

### v1

@@ 1 Introduction
Modern machine learning (ML) techniques are quickly becoming ubiquitous in the natural sciences due to their excellent data processing capabilities and ability to find excellent interpolations from high-dimensional data. The situation is not too different in various branches of physics (see ref.RevModPhys.91.045002 for a recent review), where they are being employed to enhance traditional ways of studying physical systems ranging from finding the ground state of many-body quantum systems to predicting cosmological parameters. The situation is similar in collider phenomenology. Particularly relevant for the huge data that the Large Hadron Collider (LHC) will generate in the high-luminosity runs Elmsheuser:2020omj ; HEPSoftwareFoundation:2020daq ; Held:2024gwj ; Motta:2024wyo , the community is currently striving to work out the intricacies of these efficient data handling algorithms ranging from their applications to triggers Nguyen:2018ugw ; Duarte:2018ite ; Butter:2022lkf ; Bortolato:2024uqg : the stage which decides which events to store and which to discard, the study of various detector modules Paganini:2017hrr ; Farrell:2018cjr ; Qasim:2019otl ; ATL-SOFT-PUB-2020-006 ; Buhmann:2020pmy ; Biscarat:2021dlj ; calochallenge ; Adelmann:2022ozp ; Liu:2023siw , phenomenological applications Baldi:2014kfa ; Dery:2017fap ; Metodiev:2017vrx ; Kasieczka:2021xcg ; ATLAS:2020iwa ; Aarrestad:2021oeb ; Hallin:2022eoq ; ATLAS:2023ixc ; Ngairangbam:2023cps ; CMS:2024lwn : which connects the recorded data to update our current understanding of the physical universe, and foundation models Finke:2023veq ; Butter:2023fov ; Vigl:2024lat ; Heinrich:2024sbg ; Birk:2024knn ; Harris:2024sra capable of generalising across several applications.
@@
We systematically review the nature of automatic feature extraction for phenomenology at the Large Hadron Collider, taking a more expositionary approach to what is already known in mainstream machine learning literature and their relations/reinterpretations specific to LHC phenomenology. One major theme of current research in this context, is the suitability of the point cloud representation in analysing high-energy collision events. Here, events are described as an unordered set of particles, characterised by their measured properties like four-momentum, charge, etc., whose number is not constant on an event-by-event basis. This is due to the non-conservation of particle multiplicity in relativistic quantum mechanics and our innate interest in how the measured particles relate to the interactions at subnuclear length scales. As the details of these interactions reside in the correlations between the measured particles, relational structures (like a graph) between the set’s elements are often employed to expose these correlations. While we touch upon graph-based representations and overview some applications, our coverage is essentially incomplete as our aim is to elucidate the interplay between the expressive power and generalisation capabilities of deep neural networks in the context of high-energy physics. For a more detailed literature review, we refer interested readers to refs. Larkoski:2017jix ; doi:10.1142/9789811234033_0007 ; Duarte:2020ngm ; Shlomi:2020gdn ; Thais:2022iok ; Belis:2023mqs ; Workman:2022ynf ; DeZoort2023 and the living review Feickert:2021ajf .
@@ 2 Universal Approximation: theory and practice
This section presents a brief account of the mathematical structure of MLPs, their expressive power captured in terms of Universal Approximation Theorems (UAT). These UATs, generally phrased as existence theorems, deal with the ability to approximate any continous function on a compact domain upto any arbitrary precision. Such generality at times impede a practical optimisation especially when the input dimensionality is very large. The expressivity is meticulously reduced based on the input representation and the training objective by biasing the network architecture to some smaller sets of functions. This preconditions are colloquially referred to as “inductive biases”. We end the section with a discussion on the need to curtail the expressivity via such biases.
@@ 2.2 Universal approximation theorems: a bird’s eye view
The range of functions that neural networks can approximate are encapsulated within the so-called universal approximation theorems (UATs)(see HORNIK1989359 ; Cybenko1989 ; LESHNO1993861 ; VOIGTLAENDER202333 for instance). These theorems investigate the class of functions that MLPs with a certain activation can approximate with finite but arbitrary number of hidden nodes and/or layers. Most of them prove that the set of functions represented by a certain class of neural networks are dense in the set of bounded continous functions in the $n$-dimensional hypercube.
The notion of dense sets in a larger superset is a generalisation of the properties of the rational numbers $\mathbb{Q}$ and real numbers $\mathbb{R}$, with $\mathbb{Q}$ being dense in $\mathbb{R}$. This means that for any $r\in\mathbb{R}$, we can find $q\in\mathbb{Q}$ such that $|r-q|<\epsilon$ for any arbitrarily small $\epsilon>0$. To generalise this notion to more intricate scenarios, one defines a metric, which measures a “distance” between two elements of the set. A dense set $\mathcal{D}$ in a superset $\mathcal{S}$, is one where for all elements $t\in\mathcal{S}$, we can find an element $a\in\mathcal{D}$ which is arbitrarily close with respect to a metric $d(s,a)<\epsilon$. For UATs, this is generally the metric induced by the supremum norm: $\sup_{\mathbf{x}\in D}|\hat{y}(\mathbf{x})|$, which is the least upper bound of the absolute value of the function $\hat{y}$ in the domain $D$. For a parametrised function $\tilde{y}(\mathbf{x})=f(\Theta,\mathbf{x})$ and a target function $g$, we have
@@
The superset of interest for universal approximation is $C(I_{n})$, the set of bounded continuous functions [^4] Although, there are larger function spaces like the space of Lebesgue integrable functions $L^{1}(I_{n})$ and possibility of UATs in them, we will restrict ourselves to $C(I_{n})$ as finding effective training strategies for discontinuous functions are extremely difficult for the large dimensionality of modern neural networks. from the closed unit hypercube $I_{n}=[0,1]^{n}$ to $\mathbb{R}$. One of the first universal approximation theorem Cybenko1989 , proves that the set of functions represented by an MLP of one hidden layer and arbitrary but finite number of nodes with sigmoid activation is dense in $C(I_{n})$ with respect to $d(\tilde{y},\hat{y})$. This means that for any given target function $\hat{y}\in C(I_{n})$, there exists a function $\tilde{y}(\mathbf{x})=f(\Theta_{0},\mathbf{x})$ parametrised by an MLP of some finite number of nodes in the hidden layer and a point $\Theta_{0}$ in the parameter space which is an arbitrarily close approximation of $\hat{y}$. As an existence theorem, it does not talk about how to find the particular point $\Theta_{0}$, nor does it restrict it to any subset of $\mathbb{R}^{N}$, $N$ being the dimension of $\mathcal{W}$ which is the number of tunable parameters in the MLP. This dimension is controlled by the number of nodes in the hidden layer and the theorems do not provide a particular strategy to find $N$ for a target level of accuracy $\epsilon$. However, given the generality of the function space which the theorem covers, the formal notion of existence of an arbitrarily close approximation to any bounded continuous function predicated the modern interest and revival of Artificial Neural Networks.
@@
Practically, an MLP with a single layer is not a very efficient function approximator and falls within the broader class of shallow machine learning algorithms scaling_in_book . This is because for highly fluctuating functions, one may need an exponentially large number of nodes compared to deeper architectures. It was known intuitively that depth increases the effectiveness of function approximation scaling_in_book which follows from close analogy to the computational complexity for width vs depth 10.5555/27031 ; 10.1007/3-540-62034-6_33 in circuits for implementing boolean operations, as well as the known ability of biological brains to learn simpler concepts first and more abstract concepts later in life. Deeper networks are able to approximate topologically more complex functions 6697897 ; pmlr-v49-eldan16 much more efficiently than shallow networks. [^5] In ref. 6697897 , it is found that the sum of Betti numbers (a measure of topological complexity) of the subset of the domain of the function with positive outputs for a single output network scales at most as a polynomial in the number of nodes for several activation functions for a shallow network while it can scale exponentially for deeper networks. Intuitively, this can be understood JMLR:v21:20-345 from the ability of the component functions in the overall compositional chain to induce topologically discontinous changes in each mapping.
@@
As the target function’s closed form expression is not known, the approximation can only be as good as the quality and the amount available data. In HEP where synthetic data can be simulated based on well-understood theoretical formulations, the amount and quality of data is seldom a problem. Yet the generality of the functions that dense architectures can approximate results in their handicap. The efficient optimisation via gradient descent algorithms only finds the best point $\Theta_{0}$ accesible from the initial point $\Theta_{i}$. However, domain knowledge from particle physics, motivate additional restrictions on the functions beyond continuity. Such restrictions can be built into the architecture which biases the functions to (strongly) always follow a particular property, or to (weakly) partially have some desired properties. We present a qualitative discussion of why such restrictions are needed in the next subsection.
@@ 2.3 Restrictions beyond continuity: no such thing as a free lunch
As we have seen hinted in the previous description, the existence of arbitrarily close approximation via a dense neural network to any given continous function does not neccesarily translate to finding the said approximation. Due to their non-constructive nature, one has to often revert to heuristics and brute force searches in the form of hyperparameter [^6] The different choices given a base architecture like the activation function, optimisation method, learning rate, the number of hidden layers, the number of nodes in these layers etc., are known as hyperparameters. optimisation coupled with some form of gradient descent optimisation to find a workable architecture and a particular set of parameters in the weight space. Although this is manageable for moderately large input dimensionality, it becomes intractable very fast with increasing number of dimensions. This is true even for deep fully-connected networks with more than two hidden layers. To counter this issue, additional informative priors are generally employed when one goes to very high dimensional input spaces which restricts the function beyond continuity. In this subsection, we discuss the motivation and the need to assume these priors.
@@
It is often enticing to think that given the universal nature of dense neural networks, an equally universal algorithm could exists that would guarantee an efficient optimisation for all distributions. However, this is known to be untrue Wolpert1992OnTC ; 6795940 ; SCHAFFER1994259 ; 585893 ; Wolpert2021 , where it was found that without any assumption about the underlying distribution, there is no such thing as an algorithm which performs the best in comparison to any other algorithm over the full range of possible functions. A particularly relevant exposition for the choice of learning algorithms and the inherent balance in its performance in all possible learning tasks can be found in ref. SCHAFFER1994259 , [^7] Also, see ref. scaling_in_book for a discussion on the quest for artificial intellligence and its interplay with priors and these no-free-lunch theorems. where a conservation law for off-training generalisation error was proved over the full range of target functions for a fixed input distribution and, finite and fixed training size. Simply put, an algorithm performing well for a set of target functions will lose this performance and be worse than a random guess in a different set of target functions so that the overall sum balances over the full space of target functions. Therefore, if one knows some properties of the target function beforehand, it is possible (or mandatory perhaps) to devise algorithms or models which favour that particular form. This motivates building restrictions in the possible functions that a non-parametric model (neural networks for our case) can appoximate.
@@
The restriction of the possible functions beyond contunuity can be done primarily in three ways: the data representation, the architecture of the network, and the loss function including the choice of regulariser and the training procedure. The first two are related although for a given data representation, there can be competing choices of architectures based on the complexity. The last strategy can be regarded as independent of the first two while depending more on the nature of finding the optimal weights for a given data representation and architecture like convex vs non-convex optimisation, or saddle point finding vs extrema finding of the loss function. Strictly speaking, depth is also one such prior which forces the learning of hierarchical abstractions of the input data relevant to the particular function that needs to be learnt at the output layer. However, with increasing input dimensionality, this alone is not enough to guarantee efficient search of the target function and one has to include domain-specific constraints into the architecture. For instance, the structure of convolutional neural networks with local connectivity, parameter sharing etc., is motivated from the fact that images tend to have local features which are relevent for (say) classification and they can be present anywhere in the image.
@@
When the architecture is structured according to domain-specific priors, there is usually an additional fully-connected module which takes the features and process them for the final goal of the training. Therefore, the actual function that is approximated (by the downstream network) may not follow the implemented priors. The initial domain-specific modules are colloqually known as feature extractors, meaning, they replace the feature-engineering stage which builds domain-specific features normally utilised as inputs to shallow machine learning methods. Although deep-learning is a subset of machine learning algorithms, their distinction from shallow methods resides in their ability to find better functions than those relying on traditional hand-engineered features. There is a long history of the usage of shallow machine learning methods DENBY1988429 ; PhysRevLett.65.1321 in collider experiments and phenomenology (see ref. Radovic2018 for a review). These methods utilise physics-motivated variable definitions constructed by processing the low-level high-dimensional data that one records at the detectors. Although, dense feed-forward neural networks with more than one-hidden layer, i.e., with no additional domain-specific designs, qualify as deep-learning models, the introduction of domain-specific design is what sets apart the modern algorithms which do not need extensively processed data. They can utilise the very high-dimensional low-level data and mostly outperform (dependent on the suitability of the priors) shallow machine learning methods.
@@
For LHC phenomenology, all three methods of restricting the target function are utilised and they have been generally successful. However, due to the inherent nature of scientific research where favourable outcomes get published, there is a selection bias which can sometimes lead one to the false belief of out-of-the-box deep learning algorithms always performing better than shallow methods acting on physics-intuitive observables devised by experts. For instance, see refs. Moore:2018lsr ; Faucett:2020vbu for a more pragmatic standpoint on Convolutional Neural Networks’ comparison to substructure variables. Nevertheless, the wealth of prior physics knowledge also stipulates one to utilise it as much as possible while not being too restrictive to cause a performance degradation. Although the performance is only estimated a posteriori, it has been found that physics-inspired networks generally help better the performance, interpretibility, and training convergence. We will discuss these priors and their relation to the point cloud representation in the next section.

### v2

@@ 1 Introduction
Modern machine learning (ML) techniques are quickly becoming ubiquitous in the natural sciences due to their excellent data processing capabilities and ability to find excellent interpolations from high-dimensional data. The situation is not too different in various branches of physics (see ref.RevModPhys.91.045002 for a recent review), where they are being employed to enhance traditional ways of studying physical systems ranging from finding the ground state of many-body quantum systems to predicting cosmological parameters. The situation is similar in collider phenomenology. Particularly relevant for the huge data that the Large Hadron Collider (LHC) will generate in the high-luminosity runs Elmsheuser:2020omj ; HEPSoftwareFoundation:2020daq ; Held:2024gwj ; Motta:2024wyo , the community is currently striving to work out the intricacies of these efficient data handling algorithms ranging from their applications to triggers Nguyen:2018ugw ; Duarte:2018ite ; Butter:2022lkf ; Bortolato:2024uqg : the stage which decides which events to store and which to discard, the study of various detector modules Paganini:2017hrr ; Farrell:2018cjr ; Qasim:2019otl ; ATL-SOFT-PUB-2020-006 ; Buhmann:2020pmy ; Biscarat:2021dlj ; calochallenge ; Adelmann:2022ozp ; Liu:2023siw , phenomenological applications Baldi:2014kfa ; Dery:2017fap ; Metodiev:2017vrx ; Kasieczka:2021xcg ; ATLAS:2020iwa ; Aarrestad:2021oeb ; Hallin:2022eoq ; ATLAS:2023ixc ; Ngairangbam:2023cps ; CMS:2024lwn : which connects the recorded data to update our current understanding of the physical universe, and foundation models Finke:2023veq ; Butter:2023fov ; Vigl:2024lat ; Heinrich:2024sbg ; Birk_2024 ; Harris:2024sra capable of generalising across several applications.
@@
We systematically review the nature of automatic feature extraction for phenomenology at the Large Hadron Collider, taking a more expositionary approach to what is already known in mainstream machine learning literature and their relations/reinterpretations specific to LHC phenomenology. One major theme of current research in this context, is the suitability of the point cloud representation in analysing high-energy collision events. In the point cloud representation, events are described as an unordered set of particles, characterised by their measured properties like four-momentum, charge, etc., whose number is not constant on an event-by-event basis. This is due to the non-conservation of particle multiplicity in relativistic quantum mechanics and our innate interest in how the measured particles relate to the interactions at subnuclear length scales. As the details of these interactions reside in the correlations between the measured particles, relational structures (like a graph) between the set’s elements are often employed to expose these correlations. While we touch upon graph-based representations and overview some applications, our coverage is essentially incomplete as our aim is to elucidate the interplay between the expressive power and generalisation capabilities of deep neural networks in the context of high-energy physics. For a more detailed literature review, we refer interested readers to refs. Larkoski:2017jix ; doi:10.1142/9789811234033_0007 ; Duarte:2020ngm ; Shlomi:2020gdn ; Thais:2022iok ; Belis:2023mqs ; Workman:2022ynf ; DeZoort2023 ; Hashemi:2023ruu ; Hashemi:2023rgo and the living review Feickert:2021ajf .
@@ 2 Universal Approximation: theory and practice
This section presents a brief account of the mathematical structure of MLPs, their expressive power captured in terms of Universal Approximation Theorems (UAT). These UATs, generally phrased as existence theorems, deal with the ability to approximate any continuous function on a compact domain up to any arbitrary precision. Such generality at times impede a practical optimisation especially when the input dimensionality is very large. The expressivity is meticulously reduced based on the input representation and the training objective by biasing the network architecture to some smaller sets of functions. This preconditions are colloquially referred to as “inductive biases”, closely related to the paradigm of inductive learning. In such scenarios, we train a model on a specific training dataset to be able to generalise to samples not present during the training, assuming that the new samples follow the same relationship between the input and the target variable. The expected error on the unseen samples is called the generalisation error of the model and it is estimated with the help of the validation dataset. Highly expressive models generally result in very good training error with poor generalisation and frequently overfit the training data. When the architecture is suitable for the particular task at hand, it results in a better generalisation error and hence can reach lower levels of validation loss without overfitting to the training dataset. We end the section with a discussion on the need to curtail the expressivity via such biases.
@@ 2.2 Universal approximation theorems: a bird’s eye view
The range of functions that neural networks can approximate are encapsulated within the so-called universal approximation theorems (UATs)(see HORNIK1989359 ; Cybenko1989 ; LESHNO1993861 ; VOIGTLAENDER202333 for instance). These theorems investigate the class of functions that MLPs with a certain activation can approximate with finite but arbitrary number of hidden nodes and/or layers. Most of them prove that the set of functions represented by a certain class of neural networks are dense in the set of bounded continuous functions in the $n$-dimensional hypercube.
The notion of dense sets in a larger superset is a generalisation of the properties of the rational numbers $\mathbb{Q}$ and real numbers $\mathbb{R}$, with $\mathbb{Q}$ being dense in $\mathbb{R}$. This means that for any $r\in\mathbb{R}$, we can find $q\in\mathbb{Q}$ such that $|r-q|<\epsilon$ for any arbitrarily small $\epsilon>0$. To generalise this notion to more intricate scenarios, one defines a metric, which measures a “distance” between two elements of the set. A dense set $\mathcal{D}$ in a superset $\mathcal{S}$, is one where for all elements $t\in\mathcal{S}$, we can find an element $a\in\mathcal{D}$ which is arbitrarily close with respect to a metric $d(s,a)<\epsilon$. For UATs, this is generally the metric induced by the supremum norm: $\sup_{\mathbf{x}\in D}|\hat{y}(\mathbf{x})|$, which is the least upper bound of the absolute value of the function $\hat{y}$ in the domain $D$. For a parametrised function $\tilde{y}(\mathbf{x})=f(\Theta,\mathbf{x})$ and a target function $\hat{y}$, we have
@@
The superset of interest for universal approximation is $C(I_{n})$, the set of bounded continuous functions [^4] Although, there are larger function spaces like the space of Lebesgue integrable functions $L_{1}(I_{n})$ and possibility of UATs in them, we will restrict ourselves to $C(I_{n})$ as finding effective training strategies for discontinuous functions are extremely difficult for the large dimensionality of modern neural networks. from the closed unit hypercube $I_{n}=[0,1]^{n}$ to $\mathbb{R}$. One of the first universal approximation theorem Cybenko1989 , proves that the set of functions represented by an MLP of one hidden layer and arbitrary but finite number of nodes with sigmoid activation is dense in $C(I_{n})$ with respect to $d(\tilde{y},\hat{y})$. This means that for any given target function $\hat{y}\in C(I_{n})$, there exists a function $\tilde{y}(\mathbf{x})=f(\Theta_{0},\mathbf{x})$ parametrised by an MLP of some finite number of nodes in the hidden layer and a point $\Theta_{0}$ in the parameter space which is an arbitrarily close approximation of $\hat{y}$. As an existence theorem, it does not talk about how to find the particular point $\Theta_{0}$, nor does it restrict it to any subset of $\mathbb{R}^{N}$, $N$ being the dimension of $\mathcal{W}$ which is the number of tunable parameters in the MLP. This dimension is controlled by the number of nodes in the hidden layer and the theorems do not provide a particular strategy to find $N$ for a target level of accuracy $\epsilon$. However, given the generality of the function space which the theorem covers, the formal notion of existence of an arbitrarily close approximation to any bounded continuous function predicated the modern interest and revival of Artificial Neural Networks.
@@
Practically, an MLP with a single layer is not a very efficient function approximator and falls within the broader class of shallow machine learning algorithms scaling_in_book . This is because for highly fluctuating functions, one may need an exponentially large number of nodes compared to deeper architectures. It was known intuitively that depth increases the effectiveness of function approximation scaling_in_book which follows from close analogy to the computational complexity for width vs depth 10.5555/27031 ; 10.1007/3-540-62034-6_33 in circuits for implementing boolean operations, as well as the known ability of biological brains to learn simpler concepts first and more abstract concepts later in life. Deeper networks are able to approximate topologically more complex functions 6697897 ; pmlr-v49-eldan16 much more efficiently than shallow networks. [^5] In ref. 6697897 , it is found that the sum of Betti numbers (a measure of topological complexity) of the subset of the domain of the function with positive outputs for a single output network scales at most as a polynomial in the number of nodes for several activation functions for a shallow network while it can scale exponentially for deeper networks. Intuitively, this can be understood JMLR:v21:20-345 from the ability of the component functions in the overall compositional chain to induce topologically discontinuous changes in each mapping.
@@
As the target function’s closed form expression is not known, the approximation can only be as good as the quality and the amount available data. In HEP where synthetic data can be simulated based on well-understood theoretical formulations, the amount and quality of data is seldom a problem. Yet the generality of the functions that dense architectures can approximate results in their handicap. The efficient optimisation via gradient descent algorithms only finds the best point $\Theta_{0}$ accessible from the initial point $\Theta_{i}$. However, domain knowledge from particle physics, motivate additional restrictions on the functions beyond continuity. Such restrictions can be built into the architecture which biases the functions to (strongly) always follow a particular property, or to (weakly) partially have some desired properties. We present a qualitative discussion of why such restrictions are needed in the next subsection.
@@ 2.3 Restrictions beyond continuity: no such thing as a free lunch
As we have seen hinted in the previous description, the existence of arbitrarily close approximation via a dense neural network to any given continuous function does not necessarily translate to finding the said approximation. Due to their non-constructive nature, one has to often revert to heuristics and brute force searches in the form of hyperparameter [^6] The different choices given a base architecture like the activation function, optimisation method, learning rate, the number of hidden layers, the number of nodes in these layers etc., are known as hyperparameters. optimisation coupled with some form of gradient descent optimisation to find a workable architecture and a particular set of parameters in the weight space. Although this is manageable for moderately large input dimensionality, it becomes intractable very fast with increasing number of dimensions. This is true even for deep fully-connected networks with more than two hidden layers. To counter this issue, additional informative priors are generally employed when one goes to very high dimensional input spaces which restricts the function beyond continuity. In this subsection, we discuss the motivation and the need to assume these priors.
@@
It is often enticing to think that given the universal nature of dense neural networks, an equally universal algorithm could exists that would guarantee an efficient optimisation for all distributions. However, this is known to be untrue Wolpert1992OnTC ; 6795940 ; SCHAFFER1994259 ; 585893 ; Wolpert2021 , where it was found that without any assumption about the underlying distribution, there is no such thing as an algorithm which performs the best in comparison to any other algorithm over the full range of possible functions. A particularly relevant exposition for the choice of learning algorithms and the inherent balance in its performance in all possible learning tasks can be found in ref. SCHAFFER1994259 , [^7] Also, see ref. scaling_in_book for a discussion on the quest for artificial intelligence and its interplay with priors and these no-free-lunch theorems. where a conservation law for off-training generalisation error was proved over the full range of target functions for a fixed input distribution and, finite and fixed training size. Simply put, an algorithm performing well for a set of target functions will lose it’s performance and be worse than a random guess in a different set of target functions so that the overall sum balances over the full space of target functions. Therefore, if one knows some properties of the target function beforehand, it is possible (or mandatory perhaps) to devise algorithms or models which favour that particular form. This motivates building restrictions in the possible functions that a non-parametric model (neural networks for our case) can approximate.
@@
The restriction of the possible functions beyond continuity can be done primarily in three ways: the data representation, the architecture of the network, and the loss function including the choice of regulariser and the training procedure. The first two are related although for a given data representation, there can be competing choices of architectures based on the complexity. The last strategy can be regarded as independent of the first two while depending more on the nature of finding the optimal weights for a given data representation and architecture like convex vs non-convex optimisation, or saddle point finding vs extrema finding of the loss function. Strictly speaking, depth is also one such prior which forces the learning of hierarchical abstractions of the input data relevant to the particular function that needs to be learnt at the output layer. However, with increasing input dimensionality, this alone is not enough to guarantee efficient search of the target function and one has to include domain-specific constraints into the architecture. For instance, the structure of convolutional neural networks with local connectivity, parameter sharing etc., is motivated from the fact that images tend to have local features which are relevant for (say) classification and they can be present anywhere in the image.
@@
When the architecture is structured according to domain-specific priors, there is usually an additional fully-connected module which takes the features and process them for the final goal of the training. Therefore, the actual function that is approximated (by the downstream network) may not follow the implemented priors. The initial domain-specific modules are colloquially known as feature extractors, meaning, they replace the feature-engineering stage which builds domain-specific features normally utilised as inputs to shallow machine learning methods. Although deep-learning is a subset of machine learning algorithms, their distinction from shallow methods resides in their ability to find better functions than those relying on traditional hand-engineered features. There is a long history of the usage of shallow machine learning methods DENBY1988429 ; PhysRevLett.65.1321 in collider experiments and phenomenology (see ref. Radovic2018 for a review). These methods utilise physics-motivated variable definitions constructed by processing the low-level high-dimensional data that one records at the detectors. Although, dense feed-forward neural networks with more than one-hidden layer, i.e., with no additional domain-specific designs, qualify as deep-learning models, the introduction of domain-specific design is what sets apart the modern algorithms which do not need extensively processed data. They can utilise the very high-dimensional low-level data and mostly outperform (dependent on the suitability of the priors) shallow machine learning methods.
@@
For LHC phenomenology, all three methods of restricting the target function are utilised and they have been generally successful. However, due to the inherent nature of scientific research where favourable outcomes get published, there is a selection bias which can sometimes lead one to the false belief of out-of-the-box deep learning algorithms always performing better than shallow methods acting on physics-intuitive observables devised by experts. For instance, see refs. Moore:2018lsr ; Faucett:2020vbu for a more pragmatic standpoint on Convolutional Neural Networks’ comparison to substructure variables. Nevertheless, the wealth of prior physics knowledge also stipulates one to utilise it as much as possible while not being too restrictive to cause a performance degradation. Although the performance is only estimated a posteriori, it has been found that physics-inspired networks generally help better the performance, interpretability, and training convergence. We will discuss these priors and their relation to the point cloud representation in the next section.